OpenStudy (anonymous):

Help ASAP! Will fan & medal : ) The Taco Bell management requires each till to start with $100. Jasmine uses this equation to determine the amount of money to deposit after the nightly shift ends. D = r – 100. What does the variable r represent? A. The amount of money made during the shift. B. The amount she is to get paid. C. The amount of money that was in the till at the start of the shift. D. The amount of the deposit.

OpenStudy (anonymous):

I have a feeling it's D but I'm not 100% sure

OpenStudy (anonymous):

what is a till? define that first.

OpenStudy (anonymous):

i would assume that D is deposit. so the choice of D is not an option?

OpenStudy (anonymous):

D. The amount of the deposit.

OpenStudy (anonymous):

i think its A!

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

the till is the cash register. she starts by putting 100 dollars in the cash register. r is the money she makes because 100 dollars is the money she just put there for change.

OpenStudy (anonymous):

then, at the end of the night the money she has made on top of the 100 dollars is hers to keep!

OpenStudy (anonymous):

Thank you! Would you mind helping me with a few others?

OpenStudy (anonymous):

i can try!

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

What is the equation of the line that has a slope of -6 and contains the coordinate (0, 7)? A. 6x + y = 7 B. -6x + y = 7 C. 7x – 6y =0 D. -7 + 6y = 12

OpenStudy (anonymous):

C. 7x – 6y =0

OpenStudy (anonymous):

A!

OpenStudy (anonymous):

y=mx+b. m is the slope, and b is the y intercept. (0,7) is the y intercept so b= 7, and -6 is the slope so m = -6.

OpenStudy (anonymous):

If y is a function of x in y = 3x, which statement is true? A. The independent variable, y, is 3 times the dependent variable, x. B. The dependent variable, y, is 3 times the independent variable, x. C. The independent variable, x, is 3 times the dependent variable, y. D. The dependent variable, x, is 3 times the independent variable, y.

OpenStudy (anonymous):

no idea about independent and dependent variables. its c or d, check online for the definition of independent and dependent variables.

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

hope i was correct!

OpenStudy (anonymous):

The "dependent variable" represents the output or effect, or is tested to see if it is the effect. The "independent variables" represent the inputs or causes, or are tested to see if they are the cause. Other variables may also be observed for various reasons.

OpenStudy (anonymous):

i think c? not sure!

OpenStudy (anonymous):

I'd say D.. Lol

OpenStudy (anonymous):

do what you like~